A balloon blown up has volume of 500ml at 5°C balloon is distended to 7/8th of its maximum capacity 1) will it burst at 30°C. 2) calculate the minimum temperature above. Which it will burst
1
Expert's answer
2019-09-09T06:49:01-0400

Volume of balloon"=500 ml"

Volume currently filled up"(V_1)=\frac{7}{8}\times 500=437.5 ml"

Current temperature"(T_1)=5 \degree C=273+5=278 K"

"1)" "T_2=30\degree C=273+30=303 K"

Using equation,"\frac{V_1}{T_1}=\frac{V_2}{T_2}"

"\frac{437.5}{278}=\frac{V_2}{303}"

"V_2=\frac{437.5 \times 303}{278}=476.84ml"

So it is less than 500 ml and the balloon will not burst.

"2)" The balloon will burst when the volume exceeds 500 ml,so the limiting case is "V_2=500 ml"

"\frac{V_1}{T_1}=\frac{V_2}{T_2}"

"\frac{437.5}{278}=\frac{500}{T_2}"

"T_2=\frac{500 \times 278}{437.5}=317.71 K"

So,the balloon will burst at temperature greater than "317.71 K \space or \space 317.71-273=44.71 \degree C"